OpenStudy (anonymous):

Two masses, each weighing 1.0 × 103 kilograms and moving with the same speed of 12.5 meters/second, are approaching each other. They have a head-on collision and bounce off away from each other. Assuming this is a perfectly elastic collision, what will be the approximate kinetic energy of the system after the collision? A. 1.6 × 105 joules B. 2.5 × 105 joules C. 1.2 × 103 joules D. 2.5 × 103 joules

OpenStudy (anonymous):

An elastic collision is an encounter between two bodies in which the total kinetic energy of the two bodies after the encounter is equal to their total kinetic energy before the encounter.

OpenStudy (anonymous):

\[1/2 m_1 v_1^2+1/2 m_2 (v_2)^2\]

OpenStudy (anonymous):

i got 1.609375x10^4

OpenStudy (anonymous):

1/2 (1.0*10^3) (12.5)^2 +1/2 (1.0*10^3) (12.5)^2= (1.0*10^3) (12.5)^2

OpenStudy (anonymous):

okay this time i got 1.5625x10^5

OpenStudy (anonymous):

that's what I got

OpenStudy (anonymous):

alright!

OpenStudy (turingtest):

or if you want to be clever, because the masses an velocities are equal\[1/2m_1v_1^2+1/2m_2v_2^2=mv^2\]

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

Energy is not conserved during collisions. Use momentum.

OpenStudy (turingtest):

it is in completely inelastic ones

OpenStudy (anonymous):

in elastic collision , it is

OpenStudy (turingtest):

sorry elastic*

OpenStudy (anonymous):

Ah yes. Overlooked elastic.